# Rethinking Weight-Averaged Model-merging

### v2

@@ Abstract
Weight-averaged model-merging has emerged as a powerful approach in deep learning, capable of enhancing model performance without fine-tuning or retraining. However, the underlying mechanisms that explain its effectiveness remain largely unexplored. In this paper, we investigate this technique from three novel perspectives to provide deeper insights into how and why weight-averaged model-merging works: (1) we examine the intrinsic patterns captured by the learning of the model weights, through the visualizations of their patterns on several datasets, showing that these weights often encode structured and interpretable patterns; (2) we investigate model ensemble merging strategies based on averaging on weights versus averaging on features, providing detailed analyses across diverse architectures and datasets; and (3) we explore the impact on model-merging prediction stability in terms of changing the parameter magnitude, revealing insights into the way of weight averaging works as regularization by showing the robustness across different parameter scales. Our findings shed light on the “black box” of weight-averaged model-merging, offering valuable insights and practical recommendations that advance the model-merging process.
@@ 1 Introduction
Model-merging can combine multiple independently trained models into a single, efficient model without introducing extra inference computations. This approach has been applied in contexts such as natural language processing [36], computer vision [14], and their sub-areas (e.g. federated learning [30], distillation [11], adversarial learning [42, 3] and large language model [20, 43]), where the ability to merge models effectively can lead to improved model performance, efficiency, and flexibility in deployment.
Despite the practical importance of model-merging, there has been limited research on the investigation of the underlying factors that make it effective, especially weight-averaged model-merging [33] (e.g. Uniform Soups and Greedy Soups) as the mainstream techniques. Without a clear understanding of the underlying mechanisms of model-merging, researchers and practitioners may lack guidance on when and how to apply it in a principled way. This research gap motivates the need to systematically explore the underlying patterns and behaviors in weight-averaged model-merging.
In this paper, we address this gap by providing a comprehensive analysis of weight-averaged model-merging from the following three key perspectives:
@@
Based on our viewpoints and findings, we conclude with useful proposals and source code that will be publicly available to provide a good reference for the model-merging research community.
@@ 2 Related Work
Figure 1: Average of all images for each class and linear classifiers visualization for each class in Cifar100 [12] dataset. The detailed experimental analysis is in Sec. 6.1.
@@
In recent years, model-merging [26, 17, 21, 4, 2, 19, 23, 35] has attracted growing interest. This approach has led to a range of methods focused on averaging or merging weights from independently trained models to enhance model performance or generalization. For example, Wortsman et al. [33] proposed an approach of averaging, or greedily averaging, the parameters from various models to yield “Model Soups” that show more accurate performance than the original models used to produce the “soup”. Further, Yadav et al. proposed Ties-Merging [34] based on parameter averaging by considering conflicts between the model weights during training, allowing for smoother parameter alignment across models for merging. Some other approaches [1, 25] address the problem from a different perspective, focusing on finding weight merging operations via the similarity analyses of model features. Yu et al. [39] combine models trained on domain-specific data to create a domain-agnostic model merging that generalizes well across diverse distributions.
Despite these advances, research into the underlying mechanisms that make model merging effective remains limited. This gap in understanding serves as the motivation for our paper. To focus on a specific scope, we aim to offer a more thorough analysis of weight-averaged model-merging than the study in [33], which serves as the foundation work for the area.
@@ 3 The Patterns Contained in Model Weights
In this section, we will explain the effectiveness of model soups, built by averaging model weights, via the analysis of their weights from the perspective of template matching. Specifically, we explain why the weights can be viewed as templates learned from the data and build connections with an inner product which is a widely explored operation in neural networks. By treating the weights as templates, we can understand that operations in weight space as efforts to manipulate feature representations.
@@ 3.1 Connections Between Inner Product and Template Matching
The weights of the model can be interpreted as templates or prototypes for the respective class activation. A couple of decades ago, Adaboost [7] adopted heuristically designed patterns such as HAAR features [29], which are simple, hand-crafted filters that capture edges, lines, and other low-level image characteristics through weighted sums and differences in pixel intensities. During training, Adaboost iteratively adjusts the weights of weak classifiers to improve their performance on misclassified examples.
In a linear classifier, the learned weight matrix serves as simplified templates representing each class within the data, as shown in Fig. 6.1. A key mechanism enabling these templates to operate effectively is the use of inner products to apply them to the input data. When an image is presented to the model, its activation is determined by the inner product between the image’s feature representation and each class’s template. The class whose template yields the highest activation is selected as the predicted class, indicating that the template matches the input image most closely. This phenomenon also holds for deeper networks, but differently from the linear classifier, weight templates in each layer are learned to detect different “important” features that contribute the most towards the objectives.
This inner product computation is essentially a similarity measurement. In this context, the averaging operation in the weight space represents a linear combination of weight matrices (shown as Fig. 2), which can be seen as Mixup [40] operation in the weight space. The linear combinations of samples in the weight space has a similar effect as in the feature space, e.g., Word2vec [18] as the arithmetic can be applied on the embedding space. This also explains why recent research [5, 41, 10, 44] leverages “model arithmetic” of multiple models in the weight space to merge models effectively.
This concept extends to model merging, where multiple models or variations of a model are combined. In such scenarios, the combined model retains the diverse templates or patterns learned from each ingredient model. The model after merging can work more effectively than individuals because it represents a broader set of learned templates, increasing the chances that for any given input, the merged model will provide a good match.
@@ 3.2 Formal Representation from the Perspective of Template Matching
For a linear classifier, each row of the weight vector $\mathbf{w}\in\mathbb{R}^{n}$ can be interpreted as a template for a particular class, and the input vector is denoted as $\mathbf{x}\in\mathbb{R}^{n}$. The inner product between $\mathbf{w}$ and $\mathbf{x}$ measures the alignment between them:
@@
where $\|\mathbf{w}\|$ and $\|\mathbf{x}\|$ are the norms of $\mathbf{w}$ and $\mathbf{x}$, and $\theta$ is the angle between the two vectors. When $\theta=0$, then $\cos\theta=1$, which means that we have $\mathbf{w}\cdot\mathbf{x}=\|\mathbf{w}\|\|\mathbf{x}\|$, resulting in a weight matrix $\mathbf{w}$ that perfectly matches the input $\mathbf{x}$, as the inner product in (1) is maximized when the template $\mathbf{w}$ is perfectly aligned with the input $\mathbf{x}$. On the other hand, when $\theta\neq 0$, then $\cos\theta<1$, which means that the template and image features become misaligned. This mechanism also applies for the linear layers within deep learning models.
@@
[backgroundcolor=white!10,rightline=true,leftline=true,topline=true,bottomline=true,roundcorner=2mm,everyline=true,nobreak=false] Proposal 1: From the analysis above and the experimental study shown in Sec. 6.1, since weight matrices contain generalizable template information, if the merged model’s performance is suboptimal, more focus should be directed toward resolving conflicts within the individual learned patterns in the weights.
@@ 4.1 Linear Model Scenarios
$\small\overline{\mathbf{h}}=\overline{\mathbf{W}}\mathbf{x}=\frac{1}{2}\left(\mathbf{W}_{1}+\mathbf{W}_{2}\right)\mathbf{x}$ (3)
@@
In the context of linear operations, the methods in (3) and (5) are equivalent and yield the same result $\overline{\mathbf{h}}=\overline{\mathbf{h}}^{\prime}$
4.2 Impact of Non-linear Activation Functions
The equivalence between averaging on weights or on features breaks when non-linear activation functions $\phi(\cdot)$, such as ReLU or Sigmoid, are introduced. In the case of averaging on weights, the non-linear activation is applied after averaging the weight matrices:
$\small\overline{\mathbf{h}}=\phi\left(\overline{\mathbf{W}}\mathbf{x}\right)=\phi\left[\frac{1}{2}\left(\mathbf{W}_{1}+\mathbf{W}_{2}\right)\mathbf{x}\right].$ (6)
Here, the non-linearity is applied to the weighted average of $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$. When averaging on features, the activation function is applied to each output feature independently:
@@ 4.2 Impact of Non-linear Activation Functions
$\small\overline{\mathbf{h}}=\frac{1}{2}\left(\mathbf{h}_{1}+\mathbf{h}_{2}\right)=\frac{1}{2}\left[\phi\left(\mathbf{W}_{1}\mathbf{x}\right)+\phi\left(\mathbf{W}_{2}\mathbf{x}\right)\right].$ (8)
Hence, the equivalency between $\overline{\mathbf{h}}$ and $\overline{\mathbf{h}}^{\prime}$ seen in the linear models will no longer hold for non-linear models. If we take ReLU for instance, each weight matrix retains its positive contribution through non-linear activation functions, while squashing to zero the negative contributions. As a result, averaging on features tends to preserve more information than averaging on weights, especially in scenarios where the transformations $\mathbf{W}_{1}\mathbf{x}$ and $\mathbf{W}_{2}\mathbf{x}$ exhibit diverse activations, such as when $\mathbf{W}_{1}\mathbf{x}$ and $\mathbf{W}_{2}\mathbf{x}$ have different signs, where their average may tend towards zero, leading to potential loss of information after the non-linear activation.
The key difference between the two approaches lies in how they interact with non-linear activation functions. Therefore, averaging on weights tends to smooth out the contributions from $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$, which can be beneficial in certain scenarios by reducing variance or providing a regularizing effect. However, this smoothing may also lead to a loss of fine-grained feature details, especially after non-linear activation, as illustrated in Eq. 6. In contrast, averaging on features allows each weight matrix to independently influence the final output after activation, also better adapting to individual data samples, preserving more variability and potentially enhancing the model’s expressive capacity. However, model ensemble will increase the computational cost linearly with the number of ensembled models; while model-merging will not. The FLOPs of models considered in the paper can be found in the supplementary materials.
@@
[backgroundcolor=white!10,rightline=true,leftline=true,topline=true,bottomline=true,roundcorner=2mm,everyline=true,nobreak=false] Proposal 2: From the analysis above and the experimental study shown in Sec. 6.2, averaging on features can generally produce stronger expressive ability than averaging on weights. However, the averaging on weights can show relatively good performance without introducing extra inference computational costs. Hence, this forms a trade-off between computation usage and model performance when deciding between model merging or model ensemble.
@@ 5 The Model Predictions and Weight Magnitudes
Figure 4: Scatter plot comparing the accuracies of uniform soup and logits ensemble over different weights magnitudes with VGG19 model on Cifar100. The values on the X and Y axis represent the accuracies. The experimental analysis is in Sec. 6.3.
In this section, we will examine how the weight-merging behaves against magnifying the candidate model weights by different magnitude factors. This helps us understand how the model-merging works as a similar way of regularization.
@@ 5.1 Analysis of Weight Magnitude After Averaging
In this section, we analyze the effect of averaging the weights of two independent models, denoted as $\mathbf{W}_{A}$ and $\mathbf{W}_{B}$, on their respective magnitudes. We define the maximum magnitude of the weights from both models as $M=\max(|\mathbf{W}_{A}|,|\mathbf{W}_{B}|)$. When we compute the averaged weights, we obtain:
$\small\overline{\mathbf{W}}=\frac{\mathbf{W}_{A}+\mathbf{W}_{B}}{2}.$ (9)
We seek to understand the relationship between the magnitude of the averaged weights $|\overline{\mathbf{W}}|$ and the maximum magnitude $M$.
@@
$\small\|\overline{\mathbf{W}}\|=\left\|\frac{\mathbf{W}_{A}+\mathbf{W}_{B}}{2}\right\|\leq\frac{\|\mathbf{W}_{A}\|+\|\mathbf{W}_{B}\|}{2},$ (10)
@@
$\small\frac{\|\mathbf{W}_{A}\|+\|\mathbf{W}_{B}\|}{2}\leq\max(\|\mathbf{W}_{A}\|,\|\mathbf{W}_{B}\|)=M.$ (11)
Thus, it follows that $\|\overline{\mathbf{W}}\|\leq M$. This analysis confirms that averaging the weights of two models results in a new weight whose magnitude’s upper bound is equal to the maximum magnitude of the original weights, but in most cases, it will be less than that maximum. Our analysis demonstrates that the operation of averaging the weights $\mathbf{W}_{A}$ and $\mathbf{W}_{B}$ typically results in a reduction of the maximum weight magnitude.
5.2 Variance Reduction in Model Weight Averaging
Consider two models, each with corresponding weight vectors $\mathbf{W}_{A}$ and $\mathbf{W}_{B}$. Let us denote the means and variances of the weight distributions for $\mathbf{W}_{A}$ and $\mathbf{W}_{B}$ as $\mathbb{E}[\mathbf{W}_{A}]=\mu_{A}$, $\text{Var}(\mathbf{W}_{A})=\sigma_{A}^{2}$ and $\mathbb{E}[\mathbf{W}_{B}]=\mu_{B}$, $\text{Var}(\mathbf{W}_{B})=\sigma_{B}^{2}$. The average of the weights of models A and B is defined as $\overline{\mathbf{W}}=(\mathbf{W}_{A}+\mathbf{W}_{B})/2$
Our goal is to compute the variance of $\overline{\mathbf{W}}$ and compare it to the variances of $\mathbf{W}_{A}$ and $\mathbf{W}_{B}$. The expectation of the averaged weights $\overline{\mathbf{W}}$ is given by:
$\small\mathbb{E}[\overline{\mathbf{W}}]=\mathbb{E}\left[\frac{\mathbf{W}_{A}+\mathbf{W}_{B}}{2}\right]=\frac{\mathbb{E}[\mathbf{W}_{A}]+\mathbb{E}[\mathbf{W}_{B}]}{2}=\frac{\mu_{A}+\mu_{B}}{2}$ (12)

### v3

@@ Abstract
Model-merging has emerged as a powerful approach in deep learning, capable of enhancing model performance without any training. However, the underlying mechanisms that explain its effectiveness remain largely unexplored. In this paper, we investigate this technique from three novel perspectives to empirically provide deeper insights into why and how weight-averaged model-merging works: (1) we examine the intrinsic patterns captured by the learning of the model weights, through the visualizations of their patterns on several datasets, showing that these weights often encode structured and interpretable patterns and that is the essential why model-merging can work; (2) we mathematically and empirically investigate model ensemble merging strategies based on averaging on weights versus averaging on features, providing detailed analyses across diverse architectures and datasets; and (3) we explore the impact on model-merging prediction stability in terms of changing the parameter magnitude, revealing insights into the way of weight averaging works as regularization by showing the robustness across different parameter scales. Our findings shed light on the “black box” of weight-averaged model-merging, offering valuable insights and practical recommendations that advance the model-merging process. The code is available at https://github.com/billhhh/Rethink-Merge.
@@ 1 Introduction
Model-merging can combine multiple independently trained models into a single, effective model without introducing extra inference computations. This family of approaches has been applied in contexts such as natural language processing [38], computer vision [16], and their sub-areas (e.g. federated learning [32], distillation [13], adversarial learning [44, 2] and large language model [22, 45]), where the ability to merge models effectively can lead to largely improved models’ performance, efficiency, and flexibility in deployment. Despite the practical importance of model-merging, there has been limited research on the investigation of the underlying factors that make it effective.
This research gap motivates the need to innovatively explore the underlying principles of why and how the family of weight-averaged model-merging methods can work. In this paper, we bridge this gap by providing a comprehensive analysis of weight-averaged model-merging from the following three key perspectives:
@@
Based on our viewpoints and findings, we conclude with useful proposals to provide a good reference for the community.
@@ 2 Related Work
Figure 1: Average of all images for each class (a) and the linear classifier visualization for each class on Cifar100 [14] dataset (b). We train a single linear classifier for all classes then wrap it back to the size of input images. The detailed experimental analysis is in Sec. 6.1.
@@
In recent years, model-merging [28, 23, 3, 21, 25, 37] has attracted growing interest. This approach has led to a range of methods focused on merging weights from independently trained models to enhance model performance or generalization. For example, Wortsman et al. [35] proposed an approach of averaging, or greedily averaging, the parameters from various models to yield “Model Soups” that show more accurate performance than the original models used to produce the “soup” but without introducing more computations into inference. Further, Yadav et al. proposed Ties-Merging [36] based on parameter averaging by considering conflicts between the model weights, allowing for smooth parameter merging across models. Some other approaches [1, 6, 12] address the problem from a different perspective, focusing on taking weight merging operations via finding a good way to interpolate on the loss basins. Zipit [27] takes one step ahead from Git Re-Basin [1] to align weights from inter-model and intra-model based on the alignment of features between merging models. Yu et al. [41] proposed a DARE model to combine models trained on domain-specific data to create a domain-agnostic model-merging that generalizes well across diverse distributions. This approach ensures that domain knowledge is preserved while enhancing the model’s ability to generalize across different tasks. Beyond model-merging, other techniques, such as mode connectivity path ensembles [8], analyze pathways in the parameter space between models. By identifying optimal modes along these paths, these models help improve the performance after merging. Fisher Merging [19] selects parameters that approximately maximize the joint likelihood of the models’ parameter posteriors.
Despite these advances, research into the underlying mechanisms that make model-merging effective remains limited. This gap in understanding serves as the motivation for our paper. To focus on a specific scope, we aim to offer a thorough analysis from mathematical and empirical perspectives of weight-averaged model-merging different from the study in [35], which serves as the foundation work for the area.
@@ 3 The Patterns Contained in Model Weights
In this section, we will explain the reason why does the weight-averaged model-merging work via the analysis of their weights from the perspective of template matching (an example is shown Fig. 1 and Fig. 2 for a linear model; for nonlinear models, visualizations can be found in the supplementary materials). Specifically, we explain the reason that the weights can be viewed as templates learned from the data and build connections with an inner product which is a widely explored operation in neural networks. By treating the weights as templates, it is understandable that operations in weight space as efforts to manipulate feature representations.
@@ 3.1 Connections Between Inner Product and Template Matching
The weights of the model can be interpreted as templates or prototypes for the respective class activations and it has a deep connection with machine learning. A couple of decades ago, Adaboost [7] contained heuristically designed patterns such as HAAR features [31], which are simple hand-crafted filters that capture edges / lines / other low-level image features through weighted sums and differences in pixel intensities. During training, Adaboost iteratively adjusts the weights of weak classifiers to improve their performance on misclassified examples.
In a linear classification, once the weights are learned, we wrap it back to have the same size with the input images. Take Cifar100 for instance, each image within the dataset is 32$\times$32$\times$3, if we flatten it and it will become a 3072$\times$1 vector; the weight matrix size is 100$\times$3072 without considering bias; in this case, each row of the weight matrix corresponds to each class to be classified and thus each row vector (1$\times$3072) can be wrapped back to 32$\times$32$\times$3 with the same size of the original images. The learned weight matrix serves as simplified templates that represent each class within the data, as shown in Fig. 6.1.
A key mechanism enabling these templates to operate effectively is the use of inner products to apply them to the input data. When an image is presented to the model, its activation is determined by the inner product between the image features and per class weight vector. For each class, which weight vector yields the highest activation is selected as the predicted class of that vector, indicating that the “template” aligns (or similar to) the input image most closely. This phenomenon also holds for deeper networks, but differently from the linear classifier, weight templates in each layer are learned to detect different features that contribute the most towards the objectives (e.g. in the lower layer, low level features, e.g. edge / line features, are tend to be learned). The averaging operation in the weight space represents as Fig. 2, which can be seen as the Mixup [42] operation in the weight space. Therefore, the linear combinations in the weight space has a similar effect as in the feature space. This also explains why recent research [4, 43, 11, 46] applies “model arithmetic” on multiple models in the weight space can achieve the effect of feature arithmetic as Word2vec [20].
Extending to model-merging, the combined model retains the diverse patterns learned from each ingredient model. The model after merging can work more effectively than individuals because it represents a broader set of learned templates, increasing the chances that for any given input the merged model provides a good match. Also, it may provide an effect of regularization by merging different sophistically learned patterns, which enhances the model robustness and generalization ability.
@@ 3.2 Formal Representation from the Perspective of Template Matching
As the inner product computation adopted in linear classification is essentially a similarity measurement, so for a linear classifier, each row of the weight vector $\mathbf{w}\in\mathbb{R}^{n}$ can be interpreted as a template for a particular class. The input vector is denoted as $\mathbf{x}\in\mathbb{R}^{n}$. The inner product between $\mathbf{w}$ and $\mathbf{x}$ measures the alignment between them:
@@
where $\|\mathbf{w}\|$ and $\|\mathbf{x}\|$ are the norms of $\mathbf{w}$ and $\mathbf{x}$, and $\theta$ is the angle between the two vectors. When $\theta=0$, then $\cos\theta=1$, which means that we have $\mathbf{w}\cdot\mathbf{x}=\|\mathbf{w}\|\|\mathbf{x}\|$, resulting in a weight matrix $\mathbf{w}$ that perfectly matches the input $\mathbf{x}$, as the inner product in (1) is maximized when the template $\mathbf{w}$ is perfectly aligned with the input $\mathbf{x}$. On the other hand, when $\theta\neq 0$, then $\cos\theta<1$, which means that the template and image features become misaligned. This mechanism also applies for the layers within deep learning models. The visualization of deep models can be found in the supplementary materials.
@@
[backgroundcolor=white!10,rightline=true,leftline=true,topline=true,bottomline=true,roundcorner=2mm,everyline=true,nobreak=false] Proposal 1: The analyses above and the experimental study in Sec. 6.1 show weight matrices contain generalizable template information. If in practice, the merged model’s performance is suboptimal, more attention should be directed toward resolving conflicts within the individual learned patterns in the weights.
@@ 4.1 Linear Model Scenarios
$\small\overline{\mathbf{h}}=\overline{\mathbf{W}}\mathbf{x}=\frac{1}{2}\left(\mathbf{W}_{1}+\mathbf{W}_{2}\right)\mathbf{x}=\frac{1}{2}\left(\mathbf{W}_{1}\mathbf{x}+\mathbf{W}_{2}\mathbf{x}\right).$ (3)
@@
In the context of linear operations, the methods in (3) and (5) are equivalent and yield the same result $\overline{\mathbf{h}}=\overline{\mathbf{h}}^{\prime}$.
4.2 Non-linear Model Scenarios
The equivalence between averaging on weights or on features breaks when non-linear activation functions $\phi(\cdot)$, such as ReLU or Sigmoid, are involved. In the case of averaging on weights, the non-linear activation is applied after averaging the weight matrices:
$\small\overline{\mathbf{h}}=\phi\left(\overline{\mathbf{W}}\mathbf{x}\right)=\phi\left[\frac{1}{2}\left(\mathbf{W}_{1}+\mathbf{W}_{2}\right)\mathbf{x}\right]=\phi\left[\frac{1}{2}(\mathbf{W}_{1}\mathbf{x}+\mathbf{W}_{2}\mathbf{x})\right].$ (6)
When averaging on features, the activation function is applied to each output feature independently:
@@ 4.2 Non-linear Model Scenarios
$\small\overline{\mathbf{h}}^{\prime}=\frac{1}{2}\left(\mathbf{h}_{1}+\mathbf{h}_{2}\right)=\frac{1}{2}\left[\phi\left(\mathbf{W}_{1}\mathbf{x}\right)+\phi\left(\mathbf{W}_{2}\mathbf{x}\right)\right].$ (8)
In such a case, the equivalency between $\overline{\mathbf{h}}$ and $\overline{\mathbf{h}}^{\prime}$ seen in the linear models may not hold for non-linear models. Take ReLU for instance, if elements of $\mathbf{W}_{1}\mathbf{x}$ and $\mathbf{W}_{2}\mathbf{x}$ exhibit totally opposite activations (same absolute values but with different signs), where their average will tend towards zero, leading to large loss of information after the non-linear activation. Therefore, averaging on weights tends to smooth out the contributions from $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$, which can be beneficial in certain scenarios by reducing variance or providing a regularizing effect. However, this smoothing may also lead to a loss while extracting fine-grained details. In contrast, averaging on features allows each weight matrix to independently influence the final output after activation. However, averaging on features (model ensemble) increases the computational costs linearly with the number of ensembled models, while averaging on weights (model-merging) does not suffer. The FLOPs of models considered in the paper can be found in the supplementary materials.
@@
[backgroundcolor=white!10,rightline=true,leftline=true,topline=true,bottomline=true,roundcorner=2mm,everyline=true,nobreak=false] Proposal 2: From the analyses above and the experimental study shown in Sec. 6.2, averaging on features can generally produce stronger expressive ability than averaging on weights. However, the averaging on weights can show promising performance without introducing extra inference computational overhead. Hence, this forms a trade-off between computation cost and model performance while deciding between model-merging or model ensembling.
@@ 5 The Model Predictions and Weight Magnitudes
Figure 4: Scatter plot comparing the classification accuracies of model merge (uniform soup) and model ensemble (logits ensemble) over different weights magnitudes with VGG19 model on Cifar100. The values on the X and Y axis represent the accuracies. The experimental analysis is in Sec. 6.3.
In this section, we will examine how the weight-merging behaves against magnifying the candidate model weights by different magnitude factors. This helps us understand how robust model-merging works.
@@ 5.1 Analysis of Weight Magnitude After Averaging
We start from analyzing the effect of averaging the weights of two independent matrices, denoted as $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$, on their respective magnitudes. We define the maximum entry of both matrices as $M=\max(\mathbf{W}_{1},\mathbf{W}_{2})=\max_{i,j}(|(\mathbf{W}_{1})_{i,j}|,|(\mathbf{W}_{2})_{i,j}|)$. When we compute the averaged weights, we obtain:
$\small\overline{\mathbf{W}}=\frac{\mathbf{W}_{1}+\mathbf{W}_{2}}{2}.$ (9)
We seek to understand the relationship between the max entry of the averaged weights $\overline{\mathbf{W}}$ and $M$.
@@
$\small\|\overline{\mathbf{W}}\|=\left\|\frac{\mathbf{W}_{1}+\mathbf{W}_{2}}{2}\right\|\leq\frac{\|\mathbf{W}_{1}\|+\|\mathbf{W}_{2}\|}{2},$ (10)
@@
$\small\frac{\|\mathbf{W}_{1}\|+\|\mathbf{W}_{2}\|}{2}\leq\max(\mathbf{W}_{1},\mathbf{W}_{2})=M.$ (11)
Thus, it follows that $\|\overline{\mathbf{W}}\|\leq M$ combining with Eq. 10. It confirms that averaging the weights results in a new weight whose entry’s upper bound is equal to the maximum entry of the original weights. Thus, the operation of averaging the weights $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$ typically results in a reduction of the maximum weight magnitude. This conclusion also extends to complex deep neural networks.
5.2 Weight Variance Reduces while Averaging
We denote the means and variances of $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$ as $\mathbb{E}(\mathbf{W}_{1})=\mu_{1}$, $\text{Var}(\mathbf{W}_{1})=\sigma_{1}^{2}$ and $\mathbb{E}(\mathbf{W}_{2})=\mu_{2}$, $\text{Var}(\mathbf{W}_{2})=\sigma_{2}^{2}$. Our goal is to compute the variance of $\overline{\mathbf{W}}$ and compare it to the variances of $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$. The expectation of the averaged weights $\overline{\mathbf{W}}$ is given by:
$\small\mathbb{E}(\overline{\mathbf{W}})=\mathbb{E}\left(\frac{\mathbf{W}_{1}+\mathbf{W}_{2}}{2}\right)=\frac{\mathbb{E}(\mathbf{W}_{1})+\mathbb{E}(\mathbf{W}_{2})}{2}=\frac{\mu_{1}+\mu_{2}}{2}$ (12)